Question 1 8 pts f(x) g(x) 12 (11 10 13 If H(x) = f(g(x)), what is H(10)? None of the choices 169 13 12 132
Answer #1

we have H(x) = f(g(x)) hence we can write,

H(x) = f(f(g(x)))

Using chain rule we can write,

H(x) = f(g(x))...(g())

H(x) = f(g(x))(2)

Hence,

H(10) = f(g(10))/(10)

we have g(10) = 12 and g'(10) = 11 hence,

H(10) = f(12)-11

we have f'(12) = 12 hence,

H(10) = 12.11

H(10) = 132